# A cutting-surface consensus approach for distributed robust optimization of multi-agent systems

### v2

@@ 3.1 Approximation Problem
[…]
The problem ($\rm{ADRCP}^{k}$) is a standard distributed optimization problem with a finite number of local inequality constraints and a global constraint set over time-varying unbalanced directed graphs. In order to solve the ($\rm{ADRCP}^{k}$), we propose a distributed projected gradient (DPG) algorithm under the UJSC assumption and prove the convergence and convergence rate of this algorithm. To show these results, we first introduce the following assumptions.
Assumption 3 (Solvability, Interior Point for ${\rm(ADRCP^{k})}$)
Let $\mathcal{F}^{k}_{i}=\{x\in X|g_{i}(x,y_{i})$ $\leq-\epsilon^{k}_{i},\forall y_{i}\in Y^{k}_{i}\}$ for any $i\in\mathcal{V}$, and $\mathcal{F}^{k}=\cap_{i=1}^{m}\mathcal{F}^{k}_{i}$. Assume that $\mathcal{F}^{k}$ is not an empty set, and there exists a point $\tilde{x}\in\mathcal{F}^{k}$ satisfying
$g_{i}(\tilde{x},y_{i})<-\epsilon^{k}_{i},\forall y_{i}\in Y_{i}^{k}\quad i=1,...,m$
Under Assumption 3, the ($\rm{ADRCP}^{k}$) has a nonempty set of optimal solutions, i.e., $\mathcal{F}^{k,*}\neq\emptyset$. Note that a nonempty optimal solution set is the sufficient and necessary condition for the convergence of the distributed subgradient projected algorithm [40].
@@
    There are self-loops in all $\mathcal{G}(t)$. Besides, there exists a constant $10<\gamma<$ 1 such that $a_{ij}(t)\geq\gamma$ if $a_{ij}(t)>0$, and $a_{ii}(t)\geq\gamma$ for $\forall i,j\in\mathcal{V}$ and $\forall t\in\mathbb{Z}$.
Assumption 4(i) says that each agent employs a convex combination of all agents’ states in the distributed optimization algorithms. Assumption 4(ii) ensures the agents’ information interacts at a nonvanishing rate. Without this assumption, information from some agents may gradually lose its impact over time, and in the limit, leading to loss of information from these agents [35].
@@
Then, we start to design an optimization algorithm for solving the ($\rm{ADRCP}^{k}$). In order to coordinate with the subsequent distributed cutting-surface consensus algorithm, we consider two main aspects in the algorithm design: One is that the algorithm can be applied to time-varying unbalanced directed networks under the UJSC assumption, same as the cutting-surface consensus algorithm, which will not affect the applicability and practical value of the cutting-surface consensus algorithm. The other is that the algorithm is capable of solving the ($\rm{ADRCP}^{k}$) with nonidentical local constraints, and the resulting solution of each agent is feasible to its local constraints. To the best of our knowledge, existing distributed optimization algorithms (e.g. [35, 40, 41, 42, 43, 44]) are unable to take both of these requirements into account (see Table 1), therefore, we propose a new distributed projected gradient (DPG) algorithm based on the following two steps:
Step (i): Proceed the following epigraphic reformulation for the ($\rm{ADRCP}^{k}$) to handle the unbalancedness of the distributed network:
$\mathop{\min}\limits_{(x,u)\in{\Theta}}$ $\sum^{m}_{i=1}\frac{\mathbf{1}^{\top}_{m}u}{m}$ (1)
$\quad\rm{s.t.}$ $f_{i}(x)-e^{\top}_{i}u\leq 0,\ i=1,2,...,m,$
$g_{i}({x},y_{i})\leq-\epsilon^{k}_{i}\quad\forall y_{i}\in Y^{k}_{i},\ i=1,2,...,m,$
$\ \ \rm{s.t.}$ $f_{i}(\theta)\leq 0,\quad\quad i=1,2,...,m$
$g_{i}(\theta)\leq 0,\quad\quad i=1,2,...,m$
where $\theta=(x,u)\in\Theta$, the local feasible set of agent $i$ is $\Omega_{i}=\left\{\theta\in\Theta|f_{i}(\theta)\leq 0\wedge g_{i}(\theta)\leq 0\right\}$, and the global feasible set is $\Omega=\cap_{i=1}^{m}\Omega_{i}$.
Step (ii): Utilize the subgradient projection algorithm in [35] to resolve (2):
$\theta_{i}(t+1)=P_{\Omega_{i}}\bigg{[}\sum_{j=1}^{m}a_{ij}(t)\theta_{j}(t)-\alpha(t)c\bigg{]}$ (3)
Remark 3 (Dealing with Unbalancedness)
[…]
Finally, we prove the convergence and convergence rate of the DPG algorithm (see Appendix). Here are the results we obtained:
Theorem 1 (Convergence of DPG Algorithm)
[…]
Theorem 2 (Convergence Rate of DPG Algorithm)
Let Assumptions 2, 3, 4 be satisfied. If $\left\{\alpha(t)\right\}$ is nonincreasing and $\alpha(t)=O({1}/{\sqrt{t}})$, then for the distributed projected gradient algorithm, we have $|f_{0}(\hat{\theta}_{i}(t))-f_{0}^{*}|=O({\ln t}/{\sqrt{t}})$, where $\hat{\theta}_{i}(t)={\sum_{r=0}^{t}\alpha(r)\theta_{i}(r)}/{\sum_{\tau=0}^{t}\alpha(\tau)}$.
Table 1: Comparisons between the proposed DPG algorithm and the existing subgradient projected algorithms
@@
3.2 Finite-Time Termination Algorithm
@@
In Section 3.1, we developed a DPG algorithm to solve the ($\rm{ADRCP}^{k}$) and proved that the sequence $\{\hat{\theta}_{i}(t)\}_{t}$ of each agent converges at a sublinear rate towards $\theta^{*}$. However, in practical applications, we cannot make the algorithm iterate infinitely to obtain the optimal solution $\theta^{*}$. Therefore, it is essential to design an effective termination algorithm such that the DPG algorithm terminates after a finite number of time slots while all agents obtain approximate optimal solutions of the ($\rm{ADRCP}^{k}$). Currently, the most commonly used termination method in distributed optimization is to set a maximum number of iterations, which is always a conservative consideration as it is affected by the network structures, the initial values of the decision variables, and the convergence rate of the recursive algorithm [37]. In [38, 39, 37], finite-time consensus algorithms were proposed, equipped with the DPG algorithm, to obtain the approximate consensus solutions in finite time slots under strongly connected directed graphs. Nevertheless, this finite-time consensus algorithm is not sufficient for all agents to obtain approximate optimal solution. Specifically, if $\|\theta_{i}(t)-\theta_{j}(t)\|\leq\epsilon$ for any $i,j\in\mathcal{V}$, there may exist the case that for any agent $i\in\mathcal{V}$ that $\|\theta_{i}(t)-\theta^{*}\|\gg\epsilon$.
@@
Therefore, based on finite-time consensus algorithm [37] and the zeroth-order stopping conditions [36], we develop a finite-time termination algorithm for convergent distributed consensus algorithms in the discrete-time setting, which ensures that all agents obtain approximate optimal consensus solutions. Moreover, our algorithm is applicable to time-varying unbalanced directed graphs under the UJSC assumption.
@@
where (4a) refers to the consensus stopping condition, while (4b) and (4c) denote the zeroth-order stopping conditions based on the states $\theta_{i}$ and function values $f_{i}(\theta_{i})$ respectively. Note that combining conditions (4b) and (4c) prevents the DPG algorithm from terminating prematurely on a very steep slope or a flat plateau.
@@
Given $\epsilon_{1}\!>\!0$, if the state $\theta_{i}$ of agent $i\in\mathcal{V}$ satisfies $\mathop{\max}_{j\in N^{in}_{i}(t)\cup\left\{i\right\}}\|\theta_{i}(t)-\theta_{j}(t)\|\leq\epsilon_{1}$, the agent $i$ is deemed to reach local $\epsilon_{1}$-consensus.
Based on the stopping criterion (4), we propose a distributed finite-time termination algorithm. Firstly, each agent sends four-bit data to its out-neighbors between the time slots $t$ and $t+1$. The data is $[h^{1}_{i}(t),e^{1}_{i}(t),e^{2}_{i}(t),e^{3}_{i}(t)]$. The first bit $h^{1}_{i}(t)$ is used to judge whether the stopping criterion (4) is satisfied which will be shown later. The second bit $e^{1}_{i}(t)$ is used to count the latest number of consecutively attaining local $\epsilon_{1}$-consensus in the time slot $t$. The remaining two bits $e^{2}_{i}(t)$ and $e^{3}_{i}(t)$ record the latest number of consecutively satisfying (4b) and (4c) for agent $i$ and its in-neighbors in the time slot $t$.
Then, each agent calculates (5)-(8) in the time slot $t+1$ according to its and its in-neighbors’ information.
$h^{1}_{i}(t+1)\!=\!\mathop{\min}\limits_{j\in N^{in}_{i}(t)\cup\left\{i\right\}}\!\left\{h^{1}_{i}(t),e^{1}_{i}(t),e^{2}_{i}(t),e^{3}_{i}(t)\right\}\!+\!1,$ (5)
@@
Let D denote the diameter of the graph $\mathcal{G}(t:t+S)$, i.e., $D={\rm diam}(\mathcal{G}(t:t+S))$. Consider the distributed finite-time termination algorithm given by (5)-(8) under Assumption 2. If there is an agent $i\in\mathcal{V}$ satisfying $h^{1}_{i}(t)\geq SD+1$, all agents meet the stopping criterion (4) in the time slot $t$.
@@
We use the proof idea in [37] to prove this proposition. Under Assumption 2, there exists a directed path $(i,i_{1}),(i_{1},i_{2}),...,(i_{q},j)$ from $i$ to $j$ with $q\leq SD-1$ for any $i,j\in\mathcal{V}$. Since there is an agent $j\in\mathcal{V}$ satisfying $h^{1}_{j}(t)\geq SD+1$ at time slot t, then at time slot $(t-1)$, it follows that $h^{1}_{i_{q}}(t-1)\geq SD$, $e^{1}_{i_{q}}(t-1)\geq SD$, $e^{2}_{i_{q}}(t-1)\geq SD$, $e^{3}_{i_{q}}(t-1)\geq SD$. Similarly, when the time slot is $(t-2)$, it holds that $h^{1}_{i_{q-1}}(t-2)\geq SD-1$, $e^{1}_{i_{q-1}}(t-2)\geq SD-1$, $e^{2}_{i_{q-1}}(t-2)\geq SD-1$, $e^{3}_{i_{q-1}}(t-2)\geq SD-1$. Repeat the same steps, it follows that $h^{1}_{i}(t-q-1)\geq SD-q\geq 1$, $e^{1}_{i}(t-q-1)\geq SD-q\geq 1$, $e^{2}_{i}(t-q-1)\geq SD-q\geq 1$, $e^{3}_{i}(t-q-1)\geq SD-q\geq 1$.
Since the agent $i$ is arbitrary, the system has already satisfied the stopping criterion (4) at time slot $(t-q-1)$ and consistently satisfies (4) afterwards. Therefore, if there exists $i\in\mathcal{V}$ such that $h^{1}_{i}(t)\geq SD+1$, the network reaches the stopping criterion (4) in the time slot $t$.
If the diameter $D$ is unknown, we can replace it with $m-1$, where $m$ is the number of the agents in the distributed network.
@@
Under Assumptions 2, 3, 4, 5, the DPG algorithm terminates finitely by equipping it with the distributed finite-time termination algorithm given by (5)-(8). Besides, each agent obtains a feasible solution satisfying consensus and zeroth-order stopping conditions to certain tolerances.
@@
It is straightforward to derive from Theorems 1, 2 and Proposition 1.
4 Algorithm Development and Analysis
Since the constraints of the (${\rm DRCP}$) are semi-infinite constraints, which means that an infinite number of inequality constraints are imposed on finite-dimensional decision variables, solving this problem is NP-hard. We cannot solve it by the existing distributed constrained algorithms (e.g., distributed projected subgradient algorithms [35]) since they can only address optimization problems with finite constraints. This section proposes a distributed cutting-surface consensus algorithm to solve the (${\rm DRCP}$), which locates feasible solutions satisfying consensus and zeroth-order stopping conditions to certain tolerances within a finite number of iterations. Besides, the main result of finite-time termination is established.
4.1 Distributed Cutting-Surface Consensus Approach
This paper refers to the algorithms of [31, 32, 33] to propose the cutting-surface consensus approach for solving the (${\rm DRCP}$), which contains three nested selection structures. The pseudocode and graphic illustration of the algorithm are shown in Algorithm 1 and Figure 1, respectively.
Algorithm 1 Distributed cutting-surface consensus algorithm for the distributed robust convex program (${\rm DRCP}$)
0:  For any $i\in\mathcal{V}$, restriction parameter $\epsilon^{0}_{i}>0$; finite or empty set $Y^{0}_{i}\subset Y_{i}$; optimal solution candidate $z_{i}^{0}$, $\|z_{i}^{0}\|=+\inf$; decision parameter $x^{0}_{i}=\mathbf{0}\in\mathbb{R}^{n}$; iteration counter $k=0$; reduction parameter $r>1$.

### v3

@@ 3.1 Approximation Problem of DRCP and Solving Method
In the subsequent distributed cutting-surface consensus algorithm, the ($\rm{ADRCP}^{\it k}$) denotes the distributed optimization problem constructed by the finite sets $Y^{k}_{i}$ and the restriction parameters $\epsilon_{i}^{k}$ of all agents at the iteration $k$. Here, $\epsilon^{k}_{i}$ represents the restriction parameter of the agent $i$ at the $k$-th iteration of the distributed cutting-surface consensus algorithm, and $Y_{i}^{k}\subset Y_{i}$ is a finite set containing values of the uncertainty parameter explored by agent $i$ within the previous $k$ iterations of this algorithm.
The problem ($\rm{ADRCP}^{\it k}$) is a standard distributed optimization problem with a finite number of local inequality constraints and a global constraint set over time-varying unbalanced directed graphs. In order to solve the ($\rm{ADRCP}^{\it k}$), we propose a distributed projected gradient (DPG) algorithm under the UJSC assumption and prove the convergence and convergence rate of this algorithm. To show these results, we first introduce the following assumptions.
Assumption 3 (Feasibility of $\rm(ADRCP^{k})$} and Slater Point)
Let $\mathcal{F}^{k}_{i}=\{x\in X|g_{i}(x,y_{i})$ $\leq-\epsilon^{k}_{i},\forall y_{i}\in Y^{k}_{i}\}$ be the local feasible set of agent $i\in\mathcal{V}$ in the ($\rm{ADRCP}^{\it k}$). Suppose that the intersection set $\mathcal{F}^{k}=\bigcap_{i=1}^{m}\mathcal{F}^{k}_{i}$ is nonempty, and there exists a point $\tilde{x}\in\mathcal{F}^{k}$ satisfying
$g_{i}(\tilde{x},y_{i})<-\epsilon^{k}_{i},\quad\forall y_{i}\in Y_{i}^{k},\quad\forall i\in\{1,\ldots,m\}.$
Under Assumption 3, the ($\rm{ADRCP}^{\it k}$) has a nonempty optimal solution set, i.e., $\mathcal{F}^{k,*}\neq\emptyset$.
@@
    There are self-loops in all $\mathcal{G}(t)$. Additionally, there exists a constant $0<\gamma<1$ such that $a_{ij}(t)\geq\gamma$ if $a_{ij}(t)>0$, and $a_{ii}(t)\geq\gamma$ for $\forall i,j\in\mathcal{V}$ and $\forall t\in\mathbb{Z}$.
Assumption 4(i) establishes that each agent employs a convex combination of all agents’ information in the distributed optimization algorithms. Assumption 4(ii) states that all agents are assigned a uniform, non-zero lower bound on all weights corresponding to interactions with agents providing information, thus ensuring that the information from each agent influences the information of every other agent persistently in time [31].
@@
Then, we start to design an optimization algorithm for solving the ($\rm{ADRCP}^{\it k}$). In order to coordinate with the subsequent distributed cutting-surface consensus algorithm, we consider two main aspects in the algorithm design. One is that the algorithm can be applied to time-varying unbalanced directed networks under the UJSC assumption, same as the cutting-surface consensus algorithm, which will not affect the applicability and practical value of the cutting-surface consensus algorithm. The other is that the algorithm is capable of solving the ($\rm{ADRCP}^{\it k}$) with nonidentical local constraints, and the resulting solution of each agent is feasible to its constraints. To the best of our knowledge, existing distributed optimization algorithms are unable to take both of these requirements into account (see Remark 3 and Table 1 for details).
Table 1: Comparisons between the proposed DPG algorithm and the existing (sub)gradient projected algorithms
@@
Remark 3 (Dealing with Unbalance)
The techniques for handling unbalance have been extensively studied. In the seminal work, [41] proposed a reweighting technique to address fixed unbalanced networks, but the Perron vectors of the graph and global information were required. [33] developed an additional adaptive algorithm that estimates the Perron vectors, and then employs a reweighting technique without the need for global information. Nonetheless, this algorithm cannot be applied to time-varying graphs. [42, 43] proposed a push-sum consensus protocol to handle the unbalance of time-varying graphs, but this algorithm is complicated as it involves multiple nonlinear iterations. Recently, [30] used an epigraph form of a constrained optimization problem to deal with unbalance, which is applicable to UJSC networks. Inspired by [30], we construct an epigraphic reformulation of the ($\rm{ADRCP}^{\it k}$) in Step (i) of our algorithm. Different from the D-RFP algorithm proposed in [30], we adopt the Euclidean projections to ensure that the agent estimates lie within their local feasible sets, ensuring the local feasibility of the solution instead of almost sure feasibility in [30].
Therefore, we propose a new distributed projected gradient (DPG) algorithm based on the following two steps:
Step (i): Proceed the following epigraphic reformulation for the ($\rm{ADRCP}^{\it k}$) to handle the unbalance of the distributed network:
$\hskip-14.22636pt\begin{aligned} &\mathop{\min}\limits_{(x,u)\in{\Theta}}&&\sum^{m}_{i=1}\frac{\mathbf{1}^{\top}_{m}u}{m}\\ &\quad\rm{s.t.}&&f_{i}(x)-e^{\top}_{i}u\leq 0,\ i=1,2,...,m,\\ &&&g_{i}({x},y_{i})\leq-\epsilon^{k}_{i},\ \forall y_{i}\in Y^{k}_{i},\ i=1,2,...,m,\end{aligned}$ (1)
$\ \ \rm{s.t.}$ $f_{i}(\theta)\leq 0,\quad\quad i=1,2,.,m$
$g_{i}(\theta)\leq 0,\quad\quad i=1,2,.,m$
where $\theta=(x,u)\in\Theta$, the local feasible set of agent $i$ is $\Omega_{i}=\left\{\theta\in\Theta|f_{i}(\theta)\leq 0\wedge g_{i}(\theta)\leq 0\right\}$, and the global feasible set is $\Omega=\bigcap_{i=1}^{m}\Omega_{i}$.
Step (ii): Referring to the subgradient projection algorithm proposed in [31], each agent $i$ iteratively updates its local estimate $\theta_{i}(t+1)$ of the optimal solution to problem (2) by
$\theta_{i}(t+1)=P_{\Omega_{i}}{\bigg[\sum_{j\in N^{in}_{i}(t)}a_{ij}(t)\theta_{j}(t)}-\alpha(t)c\bigg].$ (3)
Remark 4
In the DPG algorithm, Euclidean projection is required for each iteration. The computational complexity of the projection is related to the projected set. If the projected set has a relatively simple structure, such as a half-space or an interval, the projected points can be obtained directly. Otherwise, a sub-optimization problem as shown below needs to be solved for each agent $i\in\mathcal{V}$ to find the projected point.
$\hskip-31.29802pt\begin{aligned} &\mathop{\min}\limits_{\theta_{i}(t+1)\in{\Theta}}&&\|\theta_{i}(t+1)-\kappa_{i}(t)\|\\ &\quad\ \rm{s.t.}&&f_{i}(\theta_{i}(t+1))\leq 0\\ &&&g_{i}(\theta_{i}(t+1))\leq 0,\end{aligned}$
where $\kappa_{i}(t)=\sum_{j\in N^{in}_{i}(t)}a_{ij}(t)\theta_{j}(t)-\alpha(t)c$.
Finally, we prove the convergence and convergence rate of the DPG algorithm (see Appendix). Here are the results:
Theorem 1 (Convergence of DPG Algorithm)
Let $\theta^{*}$ be an optimal solution of the ($\rm{ADRCP}^{\it k}$), and $\theta^{*}$ belongs to the optimal solution set $\Omega^{*}$. Under Assumptions 25, the distributed projected gradient algorithm reaches convergence to an optimal consensus solution, i.e., $\exists\theta^{*}\in\Omega^{*}:\ \lim_{t\rightarrow\infty}\theta_{i}(t)=\theta^{*},\ \forall i\in\mathcal{V}$.
Theorem 2 (Convergence Rate of DPG Algorithm)
Let Assumptions 24 be satisfied. If $\left\{\alpha(t)\right\}$ is nonincreasing and $\alpha(t)=O({1}/{\sqrt{t}})$, then for the distributed projected gradient algorithm, we have $|f_{0}(\hat{\theta}_{i}(t))-f_{0}^{*}|=O({\ln t}/{\sqrt{t}})$, where $\hat{\theta}_{i}(t)={\sum_{r=0}^{t}\alpha(r)\theta_{i}(r)}/{\sum_{\tau=0}^{t}\alpha(\tau)}$.
3.2 Distributed Finite-Time Termination Algorithm
@@
In Section 3.1, we developed a DPG algorithm to solve the ($\rm{ADRCP}^{\it k}$) and proved that the sequence $\{\hat{\theta}_{i}(t)\}_{t}$ of each agent converges at a sublinear rate towards $\theta^{*}$. However, in practical applications, we cannot make the algorithm iterate infinitely to obtain the optimal solution $\theta^{*}$. Therefore, it is essential to design an effective termination algorithm such that the DPG algorithm terminates after a finite number of time slots while all agents obtain approximate optimal solutions of the ($\rm{ADRCP}^{\it k}$). Currently, the most commonly used termination method in distributed optimization is to set maximum number of iterations, which is always a conservative consideration as it is affected by the network structures, the initial values of the decision variables, and the convergence rate of the recursive algorithm [38]. In [39, 40, 38], finite-time consensus algorithms were proposed, equipped with the DPG algorithm, to obtain the approximate consensus solutions in finite time under strongly connected directed graphs. Nevertheless, this finite-time consensus algorithm is not sufficient for all agents to obtain approximate optimal solution. Specifically, when $\|\theta_{i}(t)-\theta_{j}(t)\|\leq\epsilon$ holds for any $i,j\in\mathcal{V}$, there may exist the case that for any agent $i\in\mathcal{V}$, $\|\theta_{i}(t)-\theta^{*}\|\gg\epsilon$.
@@
Therefore, based on finite-time consensus algorithm [38] and the zeroth-order stopping conditions [37], we develop a novel finite-time termination algorithm for distributed consensus algorithms in the discrete-time setting, which ensures that all agents obtain approximate optimal consensus solutions. Moreover, our algorithm is applicable to time-varying unbalanced directed graphs under the UJSC assumption.
@@
where $t\geq 1$. (4a) refers to the consensus stopping condition, while (4b) and (4c) denote the zeroth-order stopping conditions based on the agent estimates $\theta_{i}$ and function values $f_{i}(\theta_{i})$ respectively. Note that combining conditions (4b) and (4c) prevents the DPG algorithm from terminating prematurely on a very steep slope or a flat plateau.
@@
Given $\epsilon_{1}\!>\!0$, if the local estimate $\theta_{i}$ of agent $i\in\mathcal{V}$ satisfies $\mathop{\max}_{j\in N^{in}_{i}(t)\cup\left\{i\right\}}\|\theta_{i}(t)-\theta_{j}(t)\|\leq\epsilon_{1}$, the agent $i$ is deemed to reach local $\epsilon_{1}$-consensus.
Based on the stopping criterion (4), we propose a distributed finite-time termination algorithm. Firstly, each agent sends four-bit data to its out-neighbors between the time $t$ and $t+1$. The data is $h^{1}_{i}(t),e^{1}_{i}(t),e^{2}_{i}(t),e^{3}_{i}(t)]$. The first bit $h^{1}_{i}(t)$ is used to judge whether the stopping criterion (4) is satisfied which will be shown later. The second bit $e^{1}_{i}(t)$ is used to count the latest number of times consecutively attaining local $\epsilon_{1}$-consensus at time $t$. The remaining two bits $e^{2}_{i}(t)$ and $e^{3}_{i}(t)$ record the latest number of times consecutively satisfying (4b) and (4c) for agent $i$ and its in-neighbors at time $t$.
Then, each agent calculates (5)–(8) at time $t+1$ according to its and its in-neighbors’ information.
$h^{1}_{i}(t+1)\!=\!\mathop{\min}\limits_{j\in N^{in}_{i}(t)\cup\left\{i\right\}}\!\left\{h^{1}_{j}(t),e^{1}_{j}(t),e^{2}_{j}(t),e^{3}_{j}(t)\right\}\!+\!1,$ (5)
@@
Let $D$ denote the diameter of the graph $\mathcal{G}(t:t+S)$, i.e., $D={\rm diam}(\mathcal{G}(t:t+S))$. Consider the distributed finite-time termination algorithm given by (5)–(8) under Assumption 2. If there exists an agent $i\in\mathcal{V}$ that satisfies $h^{1}_{i}(t)\geq SD+1$, then the stopping criterion (4) is satisfied for all agents of the multi-agent system at time $t$.
@@
We use the proof idea in [38] to prove this proposition. Under Assumption 2, there exists a directed path $(j,j_{1}),(j_{1},j_{2}),...,(j_{q},i)$ from $j$ to $i$ with $q\leq SD-1$ for any $i,j\in\mathcal{V}$. Since there is an agent $i\in\mathcal{V}$ satisfying $h^{1}_{i}(t)\geq SD+1$ at time t, then at time t1t-1, it follows that $h^{1}_{j_{q}}(t-1)\geq SD$, $e^{1}_{j_{q}}(t-1)\geq SD$, $e^{2}_{j_{q}}(t-1)\geq SD$, $e^{3}_{j_{q}}(t-1)\geq SD$. Similarly, when the time is t2t-2, it holds that $h^{1}_{j_{q-1}}(t-2)\geq SD-1$, $e^{1}_{j_{q-1}}(t-2)\geq SD-1$, $e^{2}_{j_{q-1}}(t-2)\geq SD-1$, $e^{3}_{j_{q-1}}(t-2)\geq SD-1$. Repeat the same steps, it follows that $h^{1}_{j}(t-q-1)\geq SD-q\geq 1$, $e^{1}_{j}(t-q-1)\geq SD-q\geq 1$, $e^{2}_{j}(t-q-1)\geq SD-q\geq 1$, $e^{3}_{j}(t-q-1)\geq SD-q\geq 1$.
Since the agent $j$ is arbitrary, the system has already satisfied the stopping criterion (4) at time tq1t-q-1 and consistently satisfies (4) afterwards. Therefore, if there exists $i\in\mathcal{V}$ such that $h^{1}_{i}(t)\geq SD+1$, the network reaches the stopping criterion (4) at time $t$.
If the diameter $D$ is unknown, we can replace it with m1m-1, where $m$ is the number of the agents in the distributed network.
Based on the DPG algorithm and the finite-time termination algorithm, a distributed finite-time optimization method for solving ($\rm{ADRCP}^{\it k}$) in finite time is formulated. The pseudocode of the method is shown in Algorithm 1. It is worth noting that in each iteration of Algorithm 1, each agent only exchanges its local estimates $\theta_{i}(t)$, $\theta_{i}(t-1)$, the corresponding function values $f_{i}(\theta_{i}(t))$, $f_{i}(\theta_{i}(t-1))$, and the termination algorithm parameters $h^{1}_{i}(t)$, $e^{1}_{i}(t)$, $e^{2}_{i}(t)$, $e^{3}_{i}(t)$ with its neighboring agents, while the local constraint functions, local cost functions, and other private information remain undisclosed, thus protecting the privacy of each agent.
Algorithm 1 Distributed projected gradient algorithm with finite-time termination for solving the ($\rm{ADRCP}^{\it k}$)
0: For all $i\in\mathcal{V}$, $\|\theta_{i}(0)\|=+{\rm Inf}$, $\|\theta_{i}(1)\|=0$, $h^{1}_{i}(1)=0$, $e^{1}_{i}(1)=0$, $e^{2}_{i}(1)=0$, $e^{3}_{i}(1)=0$; $t=1$.
0:$S$; $D$; $\epsilon_{1}$, $\epsilon_{2}$, $\epsilon_{3}$; $\alpha(t)$.
0:
1: Send $\theta_{i}(t\!-\!1)$, $\theta_{i}(t)$, $f_{i}(\theta_{i}(t\!-\!1))$, $f_{i}(\theta_{i}(t))$, $h^{1}_{i}(t)$, $e^{1}_{i}(t)$, $e^{2}_{i}(t)$ and $e^{3}_{i}(t)$ to the out-neighbors of agent $i$.
2: Update $\theta_{i}(t+1)$ according to (3), and update $h^{1}_{i}(t+1)$, $e^{1}_{i}(t+1)$, $e^{2}_{i}(t+1)$ and $e^{3}_{i}(t+1)$ according to (5)–(8).
3: Set $t\leftarrow t+1$
4:if $h^{1}_{i}(t)\geq SD+1$ then
5:  $\bullet$ $\left[{x_{i}^{k}}^{\top},{u_{i}^{k}}^{\top}\right]^{\top}\leftarrow\theta_{i}(t)$
6:  $\bullet$ return $x^{k}_{i}$, exit
7:end if
@@
Under Assumptions 25, Algorithm 1 terminates finitely and each agent obtains a local feasible solution $x^{k}_{i}$ of ($\rm{ADRCP}^{\it k}$) satisfying consensus and zeroth-order stopping conditions to certain tolerances.
@@
According to Theorem 1 and the triangle inequality, for any $i\neq j\in\mathcal{V}$, it follows that
$\mathop{\lim}\limits_{t\rightarrow\infty}\|\theta_{i}(t)-\theta_{j}(t)\|\!\leq\!\mathop{\lim}\limits_{t\rightarrow\infty}[\|\theta_{i}(t)-\theta^{*}\|\!+\!\|\theta_{j}(t)-\theta^{*}\|]\!=\!0.$
Due to the positivity of the norm, it can be obtained that $\mathop{\lim}\limits_{t\rightarrow\infty}\|\theta_{i}(t)-\theta_{j}(t)\|=0.$ Therefore, for any $\epsilon_{1}>0$,
$\exists T_{1}:\ \|\theta_{i}(T_{1})-\theta_{j}(T_{1})\|\leq\epsilon_{1},\ \forall i,j\in\mathcal{V}.$
Similarly, we can obtain that for any $\epsilon_{2},\epsilon_{3}>0$,
$\exists T_{2}:\ \|\theta_{i}(T_{2})-\theta_{i}(T_{2}-1)\|\leq\epsilon_{2},\ \forall i\in\mathcal{V},$
$\exists T_{3}:\ |f_{i}(\theta_{i}(T_{3}))-f_{i}(\theta_{i}(T_{3}-1))|\leq\epsilon_{3},\ \forall i\in\mathcal{V}.$
Therefore, we can conclude that after at most $\max\{T_{1},T_{2},T_{3}\}$ iterations of Algorithm 1, each agent stops updating and furnishes a point $x^{k}_{i}$ satisfying consensus and zeroth-order stopping conditions within certain tolerances. In addition, since the Euclidean projection operators are adopted, $x^{k}_{i}$ lies within its local feasible region $\mathcal{F}_{i}^{k}$.
3.3 Distributed Cutting-Surface Consensus Approach for Solving The DRCP
In this subsection, we propose a distributed cutting-surface consensus approach for solving the ($\rm DRCP$}), which contains three nested selection structures. The pseudocode and graphic illustration of the algorithm are shown in Algorithm 2 and Figure 1, respectively.
Algorithm 2 Distributed cutting-surface consensus algorithm
0: For all $i\in\mathcal{V}$, restriction parameter $\epsilon^{0}_{i}>0$, finite or empty set $Y^{0}_{i}\subset Y_{i}$, optimal solution candidate $z^{0}_{i}$, $\|z_{i}^{0}\|=+{\rm Inf}$; iteration counter $k=0$; reduction parameter $r>1$.